OpenStudy (anonymous):

Please help! Find cos x if sin^2 x-1/ cos x =-1

OpenStudy (anonymous):

This will help you: https://www.wolframalpha.com/input/?i=sin^2+x-1%2F+cos+x+%3D-1

OpenStudy (solomonzelman):

\(\large\color{black}{ \sin^2 x-1=-\cos^2x }\)

OpenStudy (solomonzelman):

another thing: -cos^2x / cosx = -cosx

OpenStudy (anonymous):

Is that it or do I solve it?

OpenStudy (solomonzelman):